OpenStudy (anonymous):

Help please! Charles has $8000 to invest,part at 4% and the rest at 5%. How much does he have invested at 5% if his annual investment form both investments total is $380?

OpenStudy (mathstudent55):

Solve it like we solved the other similar problem, just with these numbers.

OpenStudy (anonymous):

x+y=800 y=8000-X so .04x+.05(8000-x)=380 .04+400-.05x=380 -.01x=-20 x=2,000 x+y=8000 2000+y=8000 y=6000 is this right?

OpenStudy (mathstudent55):

Excellent work! Correct. Now you need to answer the question. The question made no mention of x and y. You chose x and y. Answer the question: "How much does he have invested at 5%... ?" $6000 at 5%